The ratio of Hispanic to Anglo students is 4:5. If the total number of students is 986, how many of them are Anglo?
Answer by checkley77(12844)

4:5=4+5=9 elements
986/9=109.555 value of each element
5*109.555=547.778 number of Anglos.
986-547.778=438.222 number of Hispanics.
Proof:
4/5=438.222/547.778
5*438.222=4*547.778
2,191.1=2,191.1